# Gradient-Based Program Repair: Fixing Bugs in Continuous Program Spaces

### v2

@@ Abstract
Automatic program repair seeks to generate correct code from buggy programs, with most approaches searching the correct program in a discrete, symbolic space of source code tokens. This symbolic search is fundamentally limited by its inability to directly reason about program behavior. We introduce Gradient-Based Program Repair (GBPR), a new paradigm that reframes program repair as continuous optimization in a differentiable numerical program space. Our core insight is to compile symbolic programs into differentiable numerical representations, enabling search in the numerical program space directly guided by program behavior. To evaluate GBPR, we present RaspBugs, a new benchmark of 1,466 buggy symbolic RASP programs and their respective numerical representations. Our experiments demonstrate that GBPR can effectively repair buggy symbolic programs by gradient-based optimization in the numerical program space, with convincing repair trajectories. To our knowledge, we are the first to state program repair as continuous optimization in a numerical program space. Our work establishes a new direction for program repair research, bridging two rich worlds: continuous optimization and program behavior.
@@ 1 Introduction
We describe Gradient-Based Program Repair (GBPR), a novel paradigm that is founded on expressing programs as numerical representations, such as embeddings or neural network weights. With this numerical program representation associated with a loss that captures the expected behavior, GBPR repairs the bug by searching in the numerical program space. This core originality of GBPR is that it considers program behavior, the expected correct one and the buggy one, as a first-class concept in the learning pipeline, directly expressed in the loss function to be optimized. To sum up, we propose to 1) transform symbolic programs into numerical programs, 2) design loss functions that capture correct behavior, and 3) optimize numerical programs with gradient descent in the program space until a repair is found.
@@
  • Gradient-Based Program Repair, a novel paradigm for program repair that performs gradient-based program search, driven by a loss function that captures correct program behavior.
@@ 5 Experiments
Overall, our experiments over 1,466 buggy transformer programs clearly demonstrate the viability of Gradient-Based Program Repair. It is effective to use gradient optimization and an input-output specification to repair a broken symbolic program. This is a paradigm shift in the field of automatic program repair.
@@
Figure 4: Repair success rates over evaluation steps, stratified by mutation order. In GBPR, an epoch is a full forward+backward pass over the I/O samples of the training dataset; in symbolic baselines, an epoch is the symbolic evaluation of one mutated program against the same I/O samples. The idea is that all three approaches are exposed to the same amount of information per epoch, in order to drive the search. Each panel shows accuracy trajectories for GBPR, GP, and BFS (higher better). For simple bugs (orders 1-2), symbolic methods achieve higher final performance since the number of potential repairs to explore is small. For complex bugs (orders 4-5), GBPR surpasses symbolic baselines, showing how gradient-based optimization can traverse complex program spaces more efficiently than symbolic search methods.
@@
To contextualize the effectiveness of Gradient-Based Program Repair, we compare it against two baselines that repair buggy RASP programs through search in the symbolic program space, as opposed to GBPR which operates in the numerical program space.
The first baseline employs a Genetic Programming (GP) approach with an evolutionary strategy with population size $\mu=16$ and offspring generation $\lambda=16$ per generation, operating directly on RASP program source code represented as Abstract Syntax Trees (ASTs). Starting from the buggy program, the initial population is seeded with the original bug plus random single-step mutations. Each generation proceeds as follows: (1) parent selection via tournament selection with $k=3$ (i.e., we select 3 programs at random and keep the one with the highest fitness), (2) mutation of selected parents to generate offspring according the mutation operators, and (3) replacement where the top $\mu$ individuals by fitness are retained from the combined pool of parents and offspring. The fitness function is defined as accuracy on the training dataset, computed by executing each candidate program on all input-output pairs and measuring sequence-level exact matches (ignoring the BOS token). The search terminates either when achieving 100% accuracy or after evaluating 500 programs (this corresponds to $\approx$15 generations if all of them are complete). The second baseline implements exhaustive breadth-first search (BFS) that systematically explores all possible mutations paths from the buggy program.
@@
Figure 4 shows repair success rates (test accuracy $\geq 99\%$) over epochs. In GBPR, an epoch is a full forward+backward pass over the I/O samples of the training dataset; in symbolic baselines, an epoch is the symbolic evaluation of one mutated program against the same I/O samples. The idea is that all three approaches are exposed to the same amount of information per epoch, in order to drive the search. Figure 4 is stratified by mutation order (one panel per order, the top left figure is for simple bugs based on single mutations, the bottom right is for complex bugs stacking five different mutations). Each panel plots pass rate trajectories for all three methods, revealing both asymptotic dynamics and final performance.
Asymptotic behavior. First, we see that all three methods exhibit asymptotic behavior after some time. GBPR asymptotically reaches a performance level after approximately 75 training epochs, while the two symbolic baselines continue improving through their respective evaluation budgets but at slower rates. This reflects fundamentally different search dynamics. Gradient-based optimization is able to quickly make use of the training signal via gradient, combining the information of the behavior for multiple I/O samples at once. On the contrary, symbolic methods do not have this powerful joint convergence over I/O samples, but instead are able to steadily explore the search space.
@@ 6 Discussion
Limitations. Our evaluation is limited to RaspBugs, our benchmark of RASP programs; broader experimentation with other programs (e.g., with Thurnherr and Scheurer (2024)) and languages is left to future work. Additionally, GBPR can only optimize parameters within the initial model architecture obtained after Tracr compilation (see scope of Tracr in section 2). If repairing a bug requires changing the structure itself (e.g., adding a new attention head), our prototype could not repair the bug. Future work is needed on symbolic-to-numerical compilation to maximize the expressivity of the numerical program space. A further practical limitation is the reliance on large input-output datasets to define the correctness loss and drive repair (we use up to 50k pairs per program; see section 5). Real-world test suites may be smaller and incomplete, limiting applicability and increasing overfitting risk under sparse specifications.
@@ 8 Conclusion
We introduced Gradient-Based Program Repair (GBPR), casting program repair as continuous optimization. By compiling symbolic programs to differentiable numerical representations and applying gradient descent to a correctness loss, GBPR effectively repairs buggy programs. This work pioneers expression program execution semantics in a continuous, optimizable space, opening new avenues for tackling fundamental programming problems via numerical optimization.

### v3

@@ Abstract
Automatic program repair seeks to generate correct code from buggy programs, with most approaches searching the correct program in a discrete, symbolic space of source code tokens. This symbolic search is fundamentally limited by its inability to directly reason about program behavior. We introduce Gradient-Based Program Repair (GBPR), a new approach that recasts program repair as continuous optimization in a differentiable numerical program space. Our core insight is to compile symbolic programs into differentiable numerical representations, enabling search in the numerical program space directly guided by program behavior. To evaluate GBPR, we present RaspBugs, a new benchmark of 1,466 buggy symbolic RASP programs and their respective numerical representations. Our experiments demonstrate that GBPR can effectively repair buggy symbolic programs by gradient-based optimization in the numerical program space, with convincing repair trajectories. To our knowledge, we are the first to state program repair as continuous optimization in a numerical program space. Our work demonstrates the feasibility of this direction for program repair research, bridging continuous optimization and program behavior.
@@ 1 Introduction
We describe Gradient-Based Program Repair (GBPR), a novel approach that is founded on expressing programs as numerical representations, such as embeddings or neural network weights. With this numerical program representation associated with a loss that captures the expected behavior, GBPR repairs the bug by searching in the numerical program space. This core originality of GBPR is that it considers program behavior, the expected correct one and the buggy one, as a first-class concept in the learning pipeline, directly expressed in the loss function to be optimized. To sum up, we propose to 1) transform symbolic programs into numerical programs, 2) design loss functions that capture correct behavior, and 3) optimize numerical programs with gradient descent in the program space until a repair is found.
@@
  • Gradient-Based Program Repair, a novel approach for program repair that performs gradient-based program search, driven by a loss function that captures correct program behavior.
@@ 5 Experiments
Overall, our experiments over 1,466 buggy transformer programs clearly demonstrate the viability of Gradient-Based Program Repair. It is effective to use gradient optimization and an input-output specification to repair a broken symbolic program in the RASP/Tracr setting, establishing a proof-of-concept for continuous program repair.
@@
Figure 4: Repair success rates (test accuracy $\geq 99\%$) over evaluation steps, stratified by mutation order. The x-axis counts gradient steps for GBPR and programs evaluated for the symbolic baselines; a wall-clock time comparison is provided in Figure 5. GP is run with a budget of 1,000 programs per bug. Each panel shows pass-rate trajectories for GBPR, GP, and BFS (higher is better). For simple bugs (orders 1–2), symbolic methods achieve higher final performance since the number of potential repairs to explore is small. For complex bugs (orders 4–5), GBPR surpasses both symbolic baselines, showing how gradient-based optimization scales better with bug complexity.
@@
To contextualize the effectiveness of Gradient-Based Program Repair, we compare it against two baselines that operate in the symbolic program space, as opposed to GBPR which operates in the numerical program space.
The first baseline employs a Genetic Programming (GP) approach with an evolutionary strategy with population size $\mu=16$ and offspring generation $\lambda=16$ per generation, operating directly on RASP program source code represented as Abstract Syntax Trees (ASTs). Starting from the buggy program, the initial population is seeded with the original bug plus random single-step mutations. Each generation proceeds as follows: (1) parent selection via tournament selection with $k=3$ (i.e., we select 3 programs at random and keep the one with the highest fitness), (2) mutation of selected parents to generate offspring according the mutation operators, and (3) replacement where the top $\mu$ individuals by fitness are retained from the combined pool of parents and offspring. The fitness function is defined as accuracy on the training dataset, computed by executing each candidate program on all input-output pairs and measuring sequence-level exact matches (ignoring the BOS token). The search terminates either when achieving 100% accuracy or after evaluating 1,000 programs (approximately 30 complete generations). We chose this budget after verifying that GP’s progress curve plateaus well before the 1,000-program mark on the majority of bugs, indicating diminishing returns from additional evaluations. The second baseline implements exhaustive breadth-first search (BFS) that systematically explores all possible mutations paths from the buggy program.
@@
Figure 4 shows repair pass rates (test accuracy $\geq 99\%$) over evaluation steps, stratified by mutation order (one panel per order: single mutations in the top left, five simultaneous mutations in the bottom right). Each panel plots pass rate trajectories for all three methods, revealing both convergence dynamics and final performance. Figure 5 provides the corresponding wall-clock time comparison, which is the more principled measure of compute efficiency.
Asymptotic behavior. First, we see that all three methods exhibit asymptotic behavior after some time. GBPR asymptotically reaches a performance level after approximately 75 training epochs, while the two symbolic baselines continue improving through their respective evaluation budgets but at slower rates. This reflects fundamentally different search dynamics. Gradient-based optimization is able to quickly make use of the training signals via gradient, combining the information of the behavior for multiple I/O samples at once. On the contrary, symbolic methods do not have this powerful joint convergence over I/O samples, but instead are able to steadily explore the search space.
@@
Wall-Clock Time Comparison. The step-based view above does not account for the different per-step costs of the three methods. GBPR training uses GPU acceleration (1/7th A100 per run), while GP and BFS are CPU-only, so wall-clock time is a more principled but also more hardware-dependent comparison. Figure 5 shows the resulting wall-clock time curves across all six base programs.
Figure 5: Repair success rates (test accuracy $\geq 99\%$) over wall-clock time, stratified by mutation order. Each panel shows pass-rate trajectories for GBPR, GP, and BFS (higher is better). For simple bugs (orders 1–2), symbolic methods retain their advantage. For complex bugs (orders 4–5), GBPR leads both GP and BFS at every time cutoff; GBPR reaches its final performance around 1 hour, while GP and BFS continue improving before plateauing later.
The advantage of GBPR is most pronounced for complex bugs. For mutation orders 4 and 5, GBPR leads both GP and BFS at every time cutoff. GBPR reaches its final performance around 1 hour (after which all jobs have finished training via early stopping): 45.4% for order-4 bugs and 33.6% for order-5 bugs. GP and BFS continue to improve beyond 1 hour but plateau later — for order-4 bugs, GP reaches 38.1% and BFS 34.9% by $\sim$6 hours; for order-5 bugs, both plateau around 2 hours (GP: 29.1%, BFS: 26.1%). At their respective convergence points, GBPR maintains a 7–8 percentage-point advantage over the next-best method on order-4 bugs, and a 4–5 percentage-point advantage on order-5 bugs.
Takeaway Under a wall-clock time budget, GBPR strictly outperforms both GP and BFS on complex, multi-location bugs (orders 4–5), where symbolic search faces a combinatorial explosion. GBPR converges around 1 hour; symbolic methods plateau later but at lower final performance.
5.4 Effect of Specification Size.
A natural question is how sensitive GBPR is to the number of input-output examples used during repair. We conducted an ablation over $N\in\{100,1{,}000,5{,}000,10{,}000,25{,}000,40{,}000\}$ specification examples, using the same gradient-step budget (500k steps) and early-stopping mechanism as in the main experiments. Three regimes emerge clearly: below N<5, [^0] N<5{,}000, symbolic methods dominate at all mutation orders; at $N\approx 10{,}000$, GBPR becomes competitive on complex bugs (orders 4–5), matching or surpassing GP and BFS; and at $N\approx 25{,}000$40,00040{,}000, GBPR clearly surpasses both symbolic baselines on complex bugs. For simple bugs (orders 1–2), symbolic methods retain their advantage throughout the tested range, as the repair search space is small and gradient descent offers no advantage there. Full results, including a per-order breakdown and figure, are provided in Appendix C.
Takeaway GBPR is competitive on complex bugs already at $N\approx 10{,}000$ examples, well below the full specification size. For simple bugs, symbolic methods remain more data-efficient. The crossover point is cleanly connected to the mutation order.
@@ 6 Discussion
Limitations. Our evaluation is limited to RaspBugs, our benchmark of RASP programs; broader experimentation with other RASP programs (e.g., with Thurnherr and Scheurer (2024)) and languages is left to future work. GBPR can only optimize parameters within the initial model architecture obtained after Tracr compilation (see scope of Tracr in section 2). If repairing a bug requires changing the structure itself (e.g., adding a new attention head), our prototype could not repair the bug. Future work is needed on symbolic-to-numerical compilation to maximize the expressivity of the numerical program space. Another limitation is the reliance on input-output datasets to define the correctness loss. Our specification-size ablation (Appendix C) shows that below N<5, [^0] N<5{,}000 examples, symbolic methods are more effective at all mutation orders; GBPR only becomes competitive on complex bugs (orders 4–5) at $N\approx 10{,}000$. Our approach thus requires a test input generator that is good enough to explore the input space at this scale.
Compiler Assumption. The RASP/Tracr setting is deliberately designed as a controlled environment in which programs compile analytically and exactly to transformer weights, removing the gap between the symbolic and numerical representations. This is a significant simplification relative to general-purpose languages like Python, where no such analytical compiler currently exists. The key question for future work is, therefore, how to construct differentiable numerical representations for broader program families, either via learned surrogates, smooth interpretation (Chaudhuri and Solar-Lezama, 2010), or latent program spaces (Bonnet and Macfarlane, 2024; Macfarlane et al., 2026). In this paper, we make no claim that GBPR directly generalises to arbitrary languages; rather, we establish its feasibility as a proof-of-concept in a setting where the compilation assumption holds by construction, and articulate it as a research direction.
@@ 8 Conclusion
We introduced Gradient-Based Program Repair (GBPR), casting program repair as continuous optimization. By compiling symbolic programs to differentiable numerical representations and applying gradient descent to a correctness loss, GBPR effectively repairs buggy programs. This work demonstrates the feasibility of expressing program execution semantics in a continuous, optimizable space, and opens new avenues for tackling fundamental programming problems via numerical optimization.
@@ Appendix B Repairing Higher-Order Mutants
Appendix C Specification Size Ablation
This appendix presents the full results of the specification-size ablation discussed in subsection 5.4. We retrain GBPR across $N\in\{100,1{,}000,5{,}000,10{,}000,25{,}000,40{,}000\}$ input-output examples, holding the gradient-step budget fixed at 500k steps. Because shuffle_dyck only has 1,635 distinct training samples, including it at $N\geq 5{,}000$ is not possible; to keep a fixed denominator across all $N$ values we exclude it entirely from this ablation, giving 1,135 buggy programs in each condition. GP and BFS baselines are restricted to the same 1,135 jobs, each with a budget of 500 evaluated programs (matching the original ablation setup; the main experiments in section 5 use an extended budget of 1,000). The validation set is scaled proportionally to $N$: we use $\lfloor N/8\rfloor$ validation samples (e.g., 5,000 at N=40, [^0] N=40{,}000 and 12 at $N=100$). The test set is always the fixed held-out 5,000 samples from the original RaspBugs specification (as described in section 4), ensuring evaluation is comparable across all $N$ values.
Mutation order N=100 N=1,000 N=5,000 N=10,000 N=25,000 N=40,000 GP BFS
1 (simplest) 41.2 43.3 49.4 49.4 55.9 58.0 86.9 100.0
2 20.3 27.1 36.3 40.8 49.7 53.7 65.3 65.8
3 20.5 23.9 26.9 33.3 38.0 47.0 49.1 42.3
4 21.6 24.4 33.0 36.9 49.4 55.1 35.8 36.9
5 (hardest) 23.0 24.0 25.0 34.0 38.0 43.0 36.0 31.0
All 25.3 29.3 35.7 39.9 47.6 52.5 59.5 60.8
Table 3: Percentage of bugs fixed (test accuracy $\geq 99\%$) by specification size $N$ and mutation order. Bold entries indicate GBPR values where GBPR surpasses both GP and BFS; underline marks the row maximum. GBPR crosses both symbolic baselines on order-4 bugs at N=10, [^0] N=10{,}000 and on order-5 bugs at N=25, [^0] N=25{,}000. For simple bugs (orders 1–2), symbolic methods retain their advantage throughout.
Figure 8: Left: Pass rate ($\geq 99\%$ accuracy) of GBPR versus specification size NN (log scale), one curve per mutation order. Right: GBPR advantage over the best symbolic baseline ($\max(\text{GP},\text{BFS})$) per order, in percentage points. The dashed line at $y=0$ marks parity with the best symbolic method. Orders 4 and 5 cross into positive territory at N=10, [^0] N=10{,}000 and N=25, [^0] N=25{,}000 respectively; orders 1–3 remain below parity throughout the tested range.